OpenStudy (anonymous):

Which of the following statements will have the same result? Explain each step in solving each one

OpenStudy (anonymous):

f(3) when f(x)=2x+2

OpenStudy (anonymous):

f^-1(4) when f(x) = 3x-4/5

OpenStudy (anonymous):

y+10=2y+1

OpenStudy (anonymous):

@Mertsj can you help me :)

OpenStudy (mertsj):

f(3) for the first one is 8

OpenStudy (mertsj):

In the second one, f inverse is(1/3)x + 4/15 so f^(-1)(4)=4/3 + 4/15=20/15+4/15=24/15=8/5

OpenStudy (mertsj):

I'm not sure what you mean by the third one because it simplifies to y = 9

OpenStudy (anonymous):

Hmm yeah so none of them equal to the same thinggg

OpenStudy (anonymous):

and where did you get 15 from?

OpenStudy (mertsj):

You need to check out the third one. I think it might be posted incorrectly. Isn't it supposed to have an x in it?

OpenStudy (mertsj):

1/3 times 4/5 = 4/15

OpenStudy (anonymous):

no its y + 10 = 2y + 1

OpenStudy (mertsj):

Then none of them have the same result.

OpenStudy (anonymous):

and for the second dont you have to change f(x) to y and change the x and y

OpenStudy (mertsj):

yes

OpenStudy (mertsj):

y=3x-4/5 x=3y-4/5 x+4/5=3y (1/3)x+4/15=y

OpenStudy (anonymous):

@Mertsj When I explain the 2nd one can I say sub. 4 into x..cause the answer comes out to 8/5 still

OpenStudy (mertsj):

yes